# A Survey on Heterogeneous Computing Using SmartNICs and Emerging Data Processing Units (Expanded Preprint)

### v1

@@ 3.1 Challenges
Of course, DPUs and SmartNICs are not flawless solutions, with plenty of challenges of their own, regarding their adoption and usage. Here are a few of the issues one will face when working with these devices [16]:
  • Computational Power: DPUs generally have energy-efficient processors like ARM processors that are not as powerful as modern multi-core processors from Intel, AMD, and NVIDIA. The number of processing cores is also less than mainstream server devices. These design choices are motivated by the need to build energy-efficient processors in DPUs, but they increase the difficulty of applying DPUs to some problem domains.
  • Memory: DPUs often have limited on-chip memory compared to general-purpose CPUs [96]. For instance, a Bluefield-2 DPU has 16 GB of main memory [67]. This forces developers and companies using them to carefully consider how data is managed and accessed, especially for workloads with large datasets. Since DPUs also have limited memory bandwidth, high-throughput workloads can also pose a challenge. Efficient data transfer between the DPU and system memory is crucial for preventing bottlenecks; to this end, some DPUs support direct memory access (DMA) between host and DPU.
  • Programming: Developing applications and configuring SmartNICs to take advantage of their capabilities can be challenging, requiring that developers be familiar with the individual programming models and APIs of each specific SmartNIC they’re working with [9]. Furthermore, their troubleshooting capabilities are often quite limited [16].
@@
  • Power and Cooling: Energy efficient as they might be, it is still difficult to add any kind of power-hungry device to pre-existing servers in a data center with power and cooling systems that were not designed for the extra load per-rack, requiring additional considerations for any would-be mass-adopters [41].
  • Expertise: These devices are more complex to use, and will often require considerable device-specific hardware, IT, and administration expertise.
@@
  • Financial Cost: SmartNICs and DPUs, being more complex than regular NICs, are of course going to be more costly.
4 DPUs on the Market
There are currently a variety of SmartNICs on the market which can be classified as DPUs; we review a few important players briefly, here, and in even more in Table 2.
@@ 4 DPUs on the Market
NVIDIA (Mellanox) BlueField BlueField-2, 3 CPU
Marvell OCTEON, ARMADA LiquidIO III, etc. CPU
Intel IPU Mount Evans, etc. FPGA + CPU
AMD (Pensando) Alveo, Elba, Capri (several) FPGA
@@
Fungible Fl F1 DPU, etc. CPU
Table 2: A few major players in the DPU market, both vendor, product line, and processor technology [3].
NVIDIA BlueField: These DPUs are powerful, and provide an optimized infrastructure designed to accommodate a wide variety of workloads in environments such as cloud computing and data centers, achieved through the offloading, accelerating, and segregating of a vast array of workloads. The BlueField-3 is pioneering line-rate packet processing at 400Gb/s, being the first to boast this capability [3].
Marvell OCTEON: These DPUs are designed with both 5G wireless infrastructure and networking in mind, and are intended to optimize not only cloud workloads, but also machine learning inference on the edge, and other applications. They’re even developing their own in-house ML engine to work with these devices. Such hardware-based (ASIC) ML/AI accelerators have enormous potential to accelerate AI algorithms, by as much as 100x over software-based implementations [3, 61].
Intel IPU: The IPU is part of Intel’s vision for end-to-end programmable networks, and their target customers are both cloud and communication service providers, and enterprises. IPUs incorporate a mixture of ASICs, Xeon CPUs, and FPGAs, and generally have similar goals to other such devices, such as reducing infrastructure overhead, and enhancing security and isolation for both data centers and tenants [14, 15].
AMD Alveo: These devices, developed by AMD, are based on FPGAs, making it possible to use them to hardware-accelerate even custom, computationally demanding tasks such as machine learning inference, data analysis, video transcoding, and more, all while boasting up to 90x performance compared to a CPU. Given the swifter evolution of algorithms than chip design cycles, the hope is that devices such as this will be able to keep up with algorithmic changes flexibly. The Alveo is considered a pioneer in the respect that this makes it one of the first smartNICs capable of offloading literally any types of tasks, all under the same platform [3].
We also note, briefly, the importance of the AWS Nitro System and Azure’s Catapult within the evolution and space of DPU technologies [3, 97, 95, 36], but as they are somewhat more proprietary systems, we will consider them less relevant to the current research around using and improving DPUs.
Table 4 provides a list of a few company whitepapers regarding DPU devices, which the reader may find helpful.
@@ 5 Programming Languages and Frameworks
The selection of programming languages, frameworks, and development tools used for SmartNICs varies based on the distinct requirements and functionalities of the SmartNIC, and the requirements of the intended application. The decision regarding the programming languages and tools one will use is often directly influenced by the specific SmartNIC vendor and model; different SmartNICs may possess distinct capacities, and vendors often furnish custom-made unique software development kits and tools that are meticulously tailored to their hardware platforms - and not others. Furthermore, the programmability attributes of SmartNICs, such as P4 support, are undergoing continual evolution, which of course also impacts the selection.
@@
DPDK / SPDK [24, 40, 73, 30, 59, 70, 19, 62, 78, 58, 91]
MPI [72, 6, 43, 29, 86, 38, 37, 13, 82, 11]
OpenMP [88, 103, 64]
@@
Table 3: Classification of publications based on software used (frameworks, libraries, and languages). Some simply make use of the software for their work on SmartNICs, while others evaluate it in SmartNIC applications, or go so far as to modify or expand upon the software for improved use with SmartNICs. Note: Many other works may use libraries we were unaware of, and so are not included.
Although other works, such as Kfoury et al. [49], go more in-depth, we at least list a number of programming models and toolsets available for programming DPUs:
  • DOCA, and other platform-specific SDKs: DOCA is NVIDIA’s extensive software development kit (SDK), and more, for BlueField DPUs, and contains virtually all of the important tools and software packages needed for developing applications to work with their hardware, both on the device and its host, aiming to facilitate and streamline setup and development. It is the go-to for working with NVIDIA DPUs, and is even intended to improve working with AI and ML on these devices [7, 68].
@@
  • MPI: Libraries implementing the Message Passing Interface (MPI) provide general-purpose parallelism without shared memory. MPI defines interfaces for inter-process communication. Communication in MPI are basically of two types - 1) peer-to-peer communication (sending and receiving messages) and 2) collective communication (e.g., broadcasting a message). NVidia Bluefield DPUs have been used to accelerate MPI implementations for collective communication functions only [6, 72]. This is because collective communication functions like all-to-all personalized communication function require more computational work among MPI processes in order to do communication buffer management. These collective communication functions often used to accelerate scientific applications in High-Performance Computing (HPC) clusters. By leveraging the parallel processing capabilities of MPI and the offloading capabilities of DPUs, researchers can significantly enhance data processing, simulations, and scientific research in general in HPC, offering the potential for breakthroughs in fields like molecular dynamics, genomics, climate modeling, astrophysics, etc [6].
@@
    Other models and libraries extend this concept of shared memory to work over a distributed system, such as PGAS models, OpenSHMEM, and OpenSNAPI [34].
  • gRPC: Google has developed an open-source Remote Procedure Call (RPC) framework known as gRPC, which boasts high performance capabilities across widely distributed systems. gRPC’s remote procdure calls work over standard network connections with a binary protocol, and are language-agnostic. This enables applications to establish seamless communication across diverse systems and languages through a clearly defined interface. While gRPC itself is not specifically tailored for DPUs, it can be deployed alongside them to enhance particular aspects of network communication and data handling. Essentially, it becomes another communication method by which one can offload responsibilities to the SmartNIC, thus decreasing CPU utilization. It is suitable for applications needing rapid data processing and low-latency communication, particularly in data centers and networking infrastructure. Nevertheless, using DPUs and gRPC together will require customized development and configuration in order to make full use of their capabilities [34].
@@
  • Hardware Description Languages: It is important to note also that FPGAs are typically programmed very differently from CPUs and ASICs, requiring various Hardware Description Languages (HDLs), such as Verilog, VHDL, etc. [35], making the more exotic FPGA-centric smartNICs likely more difficult to program than general-purpose CPU-based DPUs, although some systems (such as Alveo Vitis HLS) provide abstractions to remove this barrier [2].
5.1 Extending Existing Tools
Current NIC software development frameworks are typically optimized for ASIC NICs, and do not take into consideration the type of flexibility inherent in off-path SmartNICs’ ability to dynamically offload a broader variety of tasks than those defined by a few ASIC accelerators [98]. Some researchers are addressing this problem by adding onto the existing languages and frameworks, creating modifications or compilers specifically for DPU applications, or to re-implement standards-based frameworks, such as MPI, with SmartNICs in mind.
MPI: In one such example, Sarkauskas et al. [72] redesign MPI’s non-blocking broadcast and all-gather collectives to make use of DPU offloading, since these primitives are difficult to synchronize efficiently. They test their designs with microbenchmarks on BlueField DPUs, reducing the execution time of the osu_ibcast microbenchmark by up to 54%, and the osu_iallgather microbenchmark by up to 43%. Additionally, Mohammadreza et al. [6] proposed a framework named ‘BluesMPI’ that demonstrates a significant enhancement in the overall execution time of the OSU Micro Benchmark when utilising the MPI_Ialltoall function, with improvements of up to 44%. The use of BluesMPI results in a notable enhancement in the overall execution duration of the P3DFFT application, with improvements of up to 30%. Turalija et al. [86] also intend to extend MPI more fully into the DPU realm with a library they are developing. Chakraborty et al [13] work towards improving MPI’s bulk-synchronous failure recovery for HPC applications, by improving upon the global-restart model, resulting in a 4x speedup compared to other solutions.
OpenMP: Usman et al. [88] introduced support for using OpenMP to treat DPUs as network co-processors for offloading, presenting a design in LLVM that supports OpenMP’s standard offloading semantics. Their proposed approach offers productivity advantages for programmers working with DPUs, compared to standard OpenMP. Similarly, the work of Zhao et al. [103] is towards flexibly allocating network resources for OpenMP, using load prediction and fuzzy control, and it very effectively improves resource utilization. Furthermore, additional compiler tools, such as OpenMP Advisor [64], may also be used to improve OpenMP’s usefulness within the realm of heterogeneous architectures; although it currently only enables code offloading to GPUs (using OpenMP and machine learning), it may be extended to other devices.
P4: Expanding P4 has also been explored. Jiarong Xing et al. [98] have created a new optimization engine for the P4 compiler which prepares its programs to dynamically rearrange network pipelines during runtime. Their experiments evidenced various successful optimizations in througput or latency, with improvements ranging from small percentages up to 49% in one case, with only a small overhead for managing the live optimization software. Their experiments were performed on Bluefield-2 and Agilio CX DPUs, as well as a software SmartNIC emulator.
These few examples alone bespeak enormous room for future work in software-level optimization for SmartNICs, and they are just the tip of the iceberg.
@@ 5.2 Creating New Tools
A few works appear to create new libraries or APIs. Of course, there are a number of other works producing a variety of code bases for more specific things, so we will only mention a few, more general communication libraries here.
@@
Some time ago, Hoefler et al. introduced ‘sPIN’ [33], a portable packet-processing acceleration programming model that offloads packet processing functions to network cards and enabling the acceleration of application and system services, including for redundant in-memory file systems. This acceleration is analogous to compute acceleration using CUDA or OpenCL. They followed more recently by introducing ‘PsPIN’ [26], an open-source implementation of the sPIN model for doing computations in the NIC. Their experiments with PsPIN boast 400 Gbit/s packet-processing speeds, with only 26 ns latencies for 64 B packets, and low power consumption.
@@
Suresh et al. [82] present additional API primitives for point-to-point and group communication patterns, for offloading any type of communication pattern to a DPU, in an attempt to combat degraded latency, bottle-necking, and limitations on communication patterns present in other state-of-the-art approaches, such as MPI and OpenSHMEM. They test their work with Alltoall micro-benchmarks, as well as a P3DFFT application.
@@
Other libraries of note, which are detailed in later sections, include AccelNet, AlNiCo, IO-TCP, PEDAL, INEC, FairNIC, iPipe, and Runway, among others [24, 53, 51, 54, 77, 30, 60, 70].
6 Application Areas and Benchmarking of DPUs
In this section, we categorize the main body of papers we’ve reviewed. Initially we identify those which are broad studies, or which perform benchmarks of DPUs, then continue on with works applying SmartNICs to various applications, by which we attempt to group them. Finally we have separated out those works connecting these devices with various forms of machine learning (ML) and artificial intelligence (AI) applications, as an area of interest. As DPUs continue to evolve, their application areas and benchmarks, and the focal points of the research community, will play a pivotal role in shaping the landscape of high-performance computing and data processing.
6.1 Broad Studies and Benchmarks

### v2

@@ 2.6 Challenges
DPUs and SmartNICs are not flawless solutions, with plenty of challenges of their own regarding their adoption and usage. Here are a few of the issues one may face when working with these devices [27]:
  • 1.
    Computational Power: DPUs generally have energy-efficient processors that are not as powerful as modern high-end multi-core server processors from Intel, AMD, and NVIDIA; often, the cores of choice are Arm cores. The number of processing cores is also less than that of mainstream server devices. These design choices are motivated by the need to build energy-efficient processors in DPUs, but they increase the difficulty of applying DPUs to some problem domains, as the CPU can easily become a performance bottleneck [1].
  • 2.
    Memory: DPUs often have limited on-chip memory compared to general-purpose CPUs. For instance, a BlueField-2 DPU has 16 GB of main memory [28]. This forces developers and companies using them to carefully consider how data is managed and accessed, especially for workloads with large datasets. Since DPUs also have limited memory bandwidth, high-throughput workloads can also pose a challenge. Efficient data transfer between the DPU and system memory is crucial for preventing bottlenecks; to this end, some DPUs support direct memory access (DMA) between host and DPU.
  • 3.
    Programming: Developing applications and configuring SmartNICs to take advantage of their capabilities can be challenging, requiring that developers be familiar with the individual programming models and APIs of each specific SmartNIC they’re working with [29]. Furthermore, their troubleshooting capabilities are often quite limited [27]. While some DPU platforms do provide debugging and emulation tools [1], these devices can still generally be difficult or slow to debug, in part because access to the DPU may be indirect.
  • 4.
@@
  • 5.
    Management: Cluster management and debugging is complex, and performance bottleneck debugging with such heterogeneous devices as DPUs can be challenging. From our observations it appears that there may be research gaps around producing or porting better tools for this and related areas.
  • 6.
    Power and Cooling: Energy efficient as they might be [30, 31, 32], it is still difficult to add any kind of power-hungry device to pre-existing servers in a data center with power and cooling systems that were not designed for the extra load per-rack, requiring additional considerations for any would-be mass-adopters [11]. As an example, Karamati et al. [30] estimated as much as 6-13% more power was required for their application, though in practice this can be difficult to measure.
  • 7.
    Expertise: DPUs are complex to use, and will often require considerable device-specific hardware, IT, and administration expertise.
  • 8.
@@
  • 9.
    Financial Cost: SmartNICs and DPUs, being more complex than regular NICs, are more costly in terms of hardware, and also in terms of power consumption.
While the research in these areas is perhaps not yet fully mature, there are a number of works which at least attempt to characterize these devices and their idiosyncrasies via benchmarking, as discussed in Section 5.
2.7 DPUs on the Market
There are currently a variety of SmartNICs on the market which can be classified as DPUs; we review a few important players briefly here, and a few more in Table 2. Different DPU platforms widely varied and nuanced hardware and accelerator capacities; for deeper critical analyses of how DPU hardware differs we refer the reader to the work by Kfoury et al. [1], and Kashyap et al. and Hu et al. [25, 33] for detailed idiosyncrasies of BlueField and other DPUs.
@@ 2.7 DPUs on the Market
NVIDIA Mellanox BlueField BlueField-2, 3 CPU
NVIDIA Converged Accel. AX800, A100X, … CPU+GPU
Marvell OCTEON, ARMADA LiquidIO III, … CPU
Intel IPU Mount Evans, … CPU+FPGA
AMD (Xilinx) Alveo (several) FPGA
AMD (Pensando) Giglio, Elba, … CPU
@@
Fungible F1, S1 CPU
MangoBoost (several) (various)
Table 2: A few major players in the DPU market, both vendor, product line, and processor technology. This information was mostly gathered from Asterfusion [34], general knowledge, and product datasheets [35].
NVIDIA BlueField: These DPUs are powerful and provide an optimized infrastructure designed to accommodate a wide variety of workloads in environments such as cloud computing and data centers, achieved through the offloading, accelerating, and segregating of a vast array of workloads. The BlueField-3 is pioneering line-rate packet processing at 400 Gb/s, being the first to boast this capability [34]. NVIDIA’s converged accelerator cards expand upon the BlueField line by adding a GPU, promising heavier in-line computing capabilities [35].
Marvell OCTEON: These DPUs are designed with both 5G wireless infrastructure and networking in mind, and are intended to optimize not only cloud workloads, but also machine learning inference on the edge, and other applications. They are developing their own in-house ML engine to work with these devices [34]. Such hardware-based (ASIC) ML/AI accelerators have a large potential to accelerate AI algorithms, by as much as 100x over software-based implementations [34, 36], as discussed in Section 6.3. LiquidIO DPUs, like others, are in use by the research community [37, 1].
Intel IPU: The IPU is part of Intel’s vision for end-to-end programmable networks, and their target customers are both cloud and communication service providers and enterprises. IPUs incorporate a mixture of ASICs, Xeon CPUs, and FPGAs, and generally have goals similar to other such devices, such as reducing infrastructure overhead and improving security and isolation for both data centers and tenants [38, 8].
AMD Alveo: These devices, developed by AMD, are based on FPGAs, making it possible to use them to hardware-acceler
-ate even custom, computationally demanding tasks such as machine learning inference, data analysis, video transcoding, and more, boasting up to 90x performance compared to a CPU. Given the swifter evolution of algorithms than chip design cycles, the hope is that devices such as this will be able to keep up with algorithmic changes flexibly, which they claim to pioneer [34].
MangoBoost: We note this more recent product line as an interesting example of highly differentiated SmartNIC-like devices designed to work together, each accelerating a different aspect of a data center, such as distributed storage, GPUs, or other operations, by offloading network management tasks to them [39]. Mangoboost collaborates with AMD, and makes use of Alveo U45N technologies within their own products.
We also note, briefly, the importance of the AWS ‘Nitro’ system and Azure’s ‘Catapult’ within the evolution and space of DPU technologies [34, 13, 40, 41], but as they are somewhat more proprietary systems, we will consider them less relevant to the current research around using and improving DPUs.
A brief list of a few company whitepapers regarding DPU devices, which the reader may find helpful, can be found in Table 5 of Section 5; these are included there because they frequently include manufacturer benchmarks.
@@ 3 Programming Languages and Frameworks
The selection of programming languages, frameworks, and development tools used for SmartNICs varies based on the distinct requirements and functionalities of the SmartNIC, and the requirements of the intended application. The decision regarding the programming languages and tools one will use is often directly influenced by the specific SmartNIC vendor and model; different SmartNICs may possess distinct capacities, and vendors often furnish custom-made, unique software development kits and tools that are meticulously tailored to their hardware platforms—and not others. Furthermore, the programmability attributes of SmartNICs, such as P4 support, are undergoing continual evolution, which also impacts the selection.
@@
DPDK / SPDK [45, 37, 26, 46, 47, 48, 49, 50, 18, 51]
MPI [52, 53, 30, 54, 42, 55, 21, 43, 56]
OpenMP [57]
@@
Table 3: Classification of publications based on software used (frameworks, libraries, and languages). These works may make use of the software, evaluate it, or expand upon it, in SmartNIC applications. Many other works may use libraries we were unaware of.
Although other works, such as that by Kfoury et al. [1], go more in-depth, we at least list a number of programming models and toolsets available for programming DPUs, and we also discuss MPI, OpenMP, and gRPC, which they do not, but which are of particular significance regarding parallel and heterogeneous computing:
  • 1.
    DOCA, and other platform-specific SDKs: DOCA is NVIDIA’s extensive software development kit (SDK), and more, for BlueField DPUs, and contains virtually all of the important tools and software packages needed for developing applications to work with their hardware, both on the device and on its host, aiming to facilitate and streamline setup and development. It is the go-to for working with NVIDIA DPUs, and is even intended to improve working with AI and ML on these devices [62, 63].
@@
  • 4.
    MPI: Libraries implementing the Message Passing Interface (MPI) provide general-purpose parallelism without shared memory, MPI itself being a standard which defines interfaces for inter-process communication. Communication in MPI is basically of two types: 1) peer-to-peer communication (sending and receiving messages), and 2) collective communication (e.g., broadcasting a message, or all-to-all patterns). These collective patterns are often necessary when accelerating scientific applications in High-Performance Computing (HPC) clusters. Combining the parallel processing of MPI with DPU offloading, researchers can enhance data processing, simulations, and scientific research in general in HPC, offering the potential for breakthroughs in fields like molecular dynamics, genomics, climate modeling, astrophysics, etc [53]. We will discuss works using or expanding MPI in Section 4.
  • 5.
@@
    Other models and libraries extend this concept of shared memory to work over a distributed system, such as PGAS models, OpenSHMEM, and OpenSNAPI [59], and works regarding OpenMP will be discussed in Section 4.
  • 6.
    gRPC: Google has developed an open-source Remote Procedure Call (RPC) framework known as gRPC, which boasts high performance capabilities across widely distributed systems. gRPC’s remote procedure calls work over standard network connections with a binary protocol, and are language-agnostic. This enables applications to establish seamless communication across diverse systems and languages through a clearly defined interface. While gRPC itself is not specifically tailored for DPUs, it can be deployed alongside them to enhance particular aspects of network communication and data handling. Essentially, it becomes another communication method by which one can offload responsibilities to the SmartNIC, thus decreasing CPU utilization. It is suitable for applications needing rapid data processing and low-latency communication, particularly in data centers and networking infrastructure. Nevertheless, using DPUs and gRPC together will require customized development and configuration in order to make full use of their capabilities [59].
  • 7.
@@
    Expanding P4 has also been explored. Xing et al. [61] have created a new optimization engine for the P4 compiler which prepares its programs to dynamically rearrange network pipelines during runtime. Their experiments evidenced various successful optimizations in throughput or latency, with improvements ranging from small percentages up to 49% in one case, with only a small overhead for managing the live optimization software. Their experiments were performed on BlueField-2 and Agilio CX DPUs, as well as a software SmartNIC emulator.
  • 8.
    Hardware Description Languages: FPGAs are typically programmed very differently from CPUs and ASICs, requiring various Hardware Description Languages (HDLs), such as Verilog, VHDL, etc. [66]. This makes the more exotic, FPGA-centric SmartNICs likely more difficult to program than general-purpose CPU-based DPUs, although some systems provide abstractions to remove this barrier, such as Alveo Vitis HLS [67] or OpenNIC shell.
@@ 3.1 Creating New Tools
A few works create new libraries or APIs. Since a good number of works produce code bases for very specific things, we will primarily mention a few, more general communication libraries here.
@@
Some time ago, Hoefler et al. introduced ‘sPIN’ [68], a portable packet-processing acceleration programming model that offloads packet processing functions to network cards and enables the acceleration of application and system services, including for redundant in-memory file systems. This acceleration is analogous to compute acceleration using CUDA or OpenCL. They followed more recently by introducing ‘PsPIN’ [69], an open-source implementation of the sPIN model for doing computations in the NIC. Their experiments with PsPIN boast 400 Gbit/s packet-processing speeds and low power consumption.
@@
Suresh et al. [21] present additional API primitives for point-to-point and group communication patterns, for offloading any type of communication pattern to a DPU, in an attempt to combat degraded latency, bottle-necking, and limitations on communication patterns present in other state-of-the-art approaches, such as MPI and OpenSHMEM. They test their work with all-to-all micro-benchmarks, as well as a P3DFFT application.
@@
Other libraries of note, which are detailed in later sections, include AlNiCo, IO-TCP, PEDAL, INEC, DORM, FairNIC, iPipe, and Runway, among others [70, 71, 72, 73, 74, 26, 24, 47].
4 Parallel and Distributed Computing on DPUs
Broadly speaking, parallelism in computing can take a number of different forms, each applicable in often very different situations. We will discuss several relevant, distinct concepts here, and how they apply relative to heterogeneous, parallel, or distributed computing with DPUs.
[57] [52, 53] [75] [54] [30] [42] [58] [55] [76] [77]
Application
     Broad Topic OpenMP MPI Infrastr. Infrastr. Simulation Simulation GIS AI/ML AI/ML Optical
     Exploration of non-infrastructure application $\bullet$ $\bullet$ $CIRCLE \CIRCLE \CIRCLE \CIRCLE \CIRCLE \CIRCLE$
[…]
Hardware and Device Usage
     Involves DPU (or exotic SmartNIC) $CIRCLE \CIRCLE \CIRCLE \CIRCLE \CIRCLE \CIRCLE \CIRCLE \CIRCLE \bullet$
[…]
Contribution
     Presents an implementation or experimentation $CIRCLE \CIRCLE \CIRCLE \CIRCLE \CIRCLE \CIRCLE \bullet$
     Includes in-depth benchmark(s) $\bullet$ $CIRCLE \CIRCLE \bullet \bullet \CIRCLE \bullet$
     Speedup/throughput improvement (rounded) - $2\times$ $4\times$ - $1\times$ - $2\times$ $\times 1×1\times$ $300\times$
     Library or framework contribution $CIRCLE \CIRCLE \bullet \bullet$
[…]
     Page count + abstract 88 $6,20$ 1515 88 1212 77 33 88 1212 2121
Discussion
[…]
     Identifies further research gaps $\bullet$ $\bullet$ $CIRCLE \CIRCLE \bullet$
[…]
Parallelism
[…]
     Exhibits distributed memory parallelism $CIRCLE \CIRCLE \CIRCLE \CIRCLE \CIRCLE \CIRCLE \CIRCLE \bullet$
     Uses multiple SmartNIC nodes $CIRCLE \CIRCLE \CIRCLE \CIRCLE \CIRCLE \CIRCLE$
[…]
     Exhibits use of internal SIMD capabilities $CIRCLE \CIRCLE$
[…]
Current NIC software development frameworks are typically optimized for ASIC NICs, and do not take into consideration the type of flexibility inherent in off-path SmartNICs’ ability to dynamically offload a broader variety of tasks than those defined by a few ASIC accelerators [61]. Similarly, many parallel computing libraries are not optimized for effective use with these powerful network devices. Some researchers are addressing these problems by adding onto the existing languages and frameworks, creating modifications or compilers specifically for DPU applications, or are re-implementing standards-based frameworks, such as MPI, with SmartNICs in mind. The works outlined below, as well as those following in the remainder of the survey, appear to indicate room for significant software-level optimizations and library-level improvements for SmartNICs. Therefore, in this section we will cite both works expanding upon MPI and OpenMP, and a few prime research examples simply applying various forms of parallelism.
Table 4 is intended to provide a useful analysis and comparison of these works’ methodologies and contributions. We posit that this table’s categorization could provide a framework for the research community to use to develop a more complete taxonomy of research works in this field. Although its scope is too narrow to make broad observations from (having only 12 papers), it can give some quick intuition about each included work’s quality or extent of research, and provides a means of finding papers with certain traits.
Shared Memory (Thread-Level) Parallelism: This is often done with the aid of libraries such as OpenMP. Because memory is shared, such parallelism is typically only across CPU cores within a single device. As most DPUs are multi-core, a large portion of the works reviewed exhibit this type of parallelism simply in their use of DPU onboard resources.
However, remote direct memory access (RDMA) and network resource allocation systems may be able to circumvent this limitation, allowing for distributed OpenMP systems to work across DPU-networked clusters, as many DPUs support RDMA or similar abstractions. Usman et al. [57] introduced support for using OpenMP to treat DPUs as network co-processors for offloading, presenting a design in LLVM that supports OpenMP’s standard offloading semantics. Their proposed approach offers productivity advantages for programmers working with DPUs, compared to standard OpenMP.
The performance results of Karamati et al. [30] and Kaymak et al. [58] demonstrate that it requires considerable engineering effort to achieve worthwhile speedups when using a DPU for applications which are not embarrassingly parallel.
Distributed Memory (Device or Process-Level) Parallelism: Upon adding the limitation that each worker process has its own dedicated, unshared memory, work is typically performed on different regions of the data concurrently. Libraries such as MPI allow us to more easily divide and coordinate this work between processes and devices, even heterogeneous devices such as between hosts and DPUs, to perform cluster computing. A number of works combine MPI and DPUs [52, 53, 30, 54, 42, 55], though any cluster computing done treating DPUs as compute nodes would also fit here.
Most interesting perhaps are the works which not only use MPI, but rather extend or optimize it for use with DPUs, especially to improve upon the collective communication functions, because these require more buffer management and synchronization work among processes, and DPUs are communication-oriented devices that could help address this frequent bottleneck. In one such example, Sarkauskas et al. [52] redesign MPI’s non-blocking broadcast and all-gather collectives to make use of DPU offloading, since these primitives are difficult to synchronize efficiently. They test their designs with microbenchmarks on BlueField DPUs, reducing the execution time of osu_ibcast and osu_iallgather by up to 54% and 43%, respectively. Additionally, Mohammadreza et al. [53] proposed a framework named ‘BluesMPI’ that demonstrates a significant enhancement in the overall execution time of the OSU Micro Benchmark when utilising the MPI_Ialltoall function, with improvements of up to 44%. The use of BluesMPI results in a notable enhancement in the overall execution duration of the P3DFFT application, with improvements of up to 30%. Turalija et al. [42] also intend to extend MPI more fully into the DPU realm with a library they are developing.
SIMD (Single Instruction Multiple Data) Parallelism: Expanding upon the prior concepts, SIMD represents a hardware-accelerated application of data-level parallelism, with many small cores performing the same computations, locked in tandem, on different data items. This requires specialized hardware, such as a GPU, TPU (tensor processing unit), vector unit of a CPU, or similar device.
While some DPUs contain SIMD-capable hardware [35], and BlueField DPU processors contain SIMD vector units, most DPUs are not targeting GPU-like use cases, instead preferring to be a means of speeding up communication between a datacenter and its host’s other hardware [75, 76], which may include GPUs.
Pipeline Parallelism: Pipelining is the concept of taking a set of instructions that has multiple steps, and instead of completing all steps before moving onto the next data item, stacking them so that any non-dependent steps are performed in parallel, often requiring specialized hardware.
However, pipelining can also be taken advantage of algorithmically across devices, such as when Karamati et al. [30] take an algorithm which is not easily parallelizable under distributed memory settings and rearrange it in order to perform communication-heavy steps on DPUs concurrently with more computational steps on host devices, achieving up to 20% speedup. Another example is when Kaymak et al. [58] perform a complex query’s ‘filter’ step on a DPU, and ‘refine’ on its host. One final example is that of Zhong et al. [77], who also modify an algorithm to effectively pipeline the sending of computations to an optical computing device.
Research Gaps: Some of these works identify areas where further research could be beneficial to the community, beyond what was performed. Some topics of note include HPC offloading using DPUs as low-power accelerators [57], more sophisticated halo exchanges [57, 30] and using ASICs for communication-intensive simulations [42], algorithmic restructurings [30], in-network runtime software [54], and the rebalancing of DPU hardware designs [30].
5 Broad Studies and Benchmarking of DPUs